A sample of Ar gas with a volume of 48.6 mL at 610 torr is compressed at constant temperature until its volume is 32.7 mL. What will be the new pressure of the sample in torr?
Answer #1

We use charles law

V1/T1=V2/T2

V1= 48.6ml

T1= 610 torr

V2= 32.7 ml

T2= 32.7*610 /48.6 =410.4 torr the new pressure